# Denoising Task Difficulty-based Curriculum for Training Diffusion Models

### v2

@@ Appendix E Details on Experimental Setups
To evaluate the performance of models, we utilized three metrics: FID [16], IS [51], and Precision/Recall [30]. Specifically, we applied FID and IS to measure sample quality, while Precision is used to assess quality further and Recall was utilized to evaluate the diversity of the generated samples in ImageNet setup. In other datasets, we employed FID to evaluate sample quality. Unless otherwise mentioned, we sampled 50K samples for evaluation. In tasks involving conditional generation, including class-conditional image generation (e.g. CIFAR-10, ImageNet) and text-to-image conversion (e.g. MS-COCO), we adapted the classifier-free guidance [19] with a guidance scale of 1.5.
@@
For training diffusion models, we utilized the AdamW optimizer [36] with a constant learning rate of 0.0001, and weight decay was not applied. The exponential moving average (EMA) to the model’s weights was used to stabilize the training and the decay ratio was set to 0.9999. The batch size was set to 256, and we augmented the training data by a horizontal flip. While the diffusion timestep $T$ was configured as 1,000 for all experiments, we trained for 100K iterations for the FFHQ dataset [24], and 400K iterations for the ImageNet dataset [7] and MS-COCO dataset [33]. The number of clusters $N$ was 20 unless otherwise specified. The maximum patience $\tau$ was varied across model sizes: it was set at 200 for DiT-S/2, DiT-B/2, and EDM, and 400 for DiT-L/2. EDM was trained using fp16, while the other models were trained using fp32. We used 8 A100 GPUs for all experiments.
@@
Figure C: Qualitative comparison between vanilla, naive curriculum, and ours on MS-COCO dataset.
@@ Appendix F Qualitative Results
Figure A presents a qualitative analysis of the performance in unconditional facial image synthesis among the vanilla, the naive curriculum approach, and our method. Our approach demonstrates superior ability in generating realistic images.
@@
In the conditional image synthesis, we exhibit the outcomes generated by the vanilla, the naive curriculum strategy, and our proposed method. Figure B clearly shows that our methodology surpasses the competing approaches in terms of quality.
@@
To further substantiate the effectiveness of our proposed technique, we conduct a qualitative comparison of the results in the text-to-image generation task among the vanilla, the naive curriculum method, and our own approach, as depicted in Fig. C.
@@ G.1 Convergence Speed across Model Size
Figure D: We observed an increase in convergence speed across various model sizes when the proposed curriculum learning approach was applied.
By leveraging the advantages of curriculum learning in diffusion training, our method offers faster convergence than vanilla training. To further investigate this aspect, we measured FID-10K through training iterations for DiT-S and DiT-L. Figure D describes the results, showing that our curriculum approach achieves faster convergence in both models. These results also support the effectiveness of our method.
@@ G.3 Qualitative Results from DiT-L/2 with 2M iterations
In Figures E-I, we present images generated by DiT-L using our curriculum training method for 2M iterations. The results demonstrate that our method produces highly realistic images.
Figure E: Uncurated 256$\times$256 DiT-L/2 samples.
@@
Figure F: Uncurated 256$\times$256 DiT-L/2 samples.
@@
Figure G: Uncurated 256$\times$256 DiT-L/2 samples.
@@
Figure H: Uncurated 256$\times$256 DiT-L/2 samples.
@@
Figure I: Uncurated 256$\times$256 DiT-L/2 samples.
@@
Appendix H Broader Impacts

### v3

@@ Appendix E Details on Experimental Setups
To evaluate the performance of models, we utilized three metrics: FID (Heusel et al., 2017), IS (Salimans et al., 2016), and Precision/Recall (Kynkäänniemi et al., 2019). Specifically, we applied FID and IS to measure sample quality, while Precision is used to assess quality further and Recall was utilized to evaluate the diversity of the generated samples in ImageNet setup. In other datasets, we employed FID to evaluate sample quality. Unless otherwise mentioned, we sampled 50K samples for evaluation. In tasks involving conditional generation, including class-conditional image generation (e.g. CIFAR-10, ImageNet) and text-to-image conversion (e.g. MS-COCO), we adapted the classifier-free guidance (Ho & Salimans, 2022) with a guidance scale of 1.5.
@@
For training diffusion models, we utilized the AdamW optimizer (Loshchilov & Hutter, 2017) with a constant learning rate of 0.0001, and weight decay was not applied. The exponential moving average (EMA) to the model’s weights was used to stabilize the training and the decay ratio was set to 0.9999. The batch size was set to 256, and we augmented the training data by a horizontal flip. While the diffusion timestep $T$ was configured as 1,000 for all experiments, we trained for 100K iterations for the FFHQ dataset (Karras et al., 2019), and 400K iterations for the ImageNet dataset (Deng et al., 2009) and MS-COCO dataset (Lin et al., 2014). The number of clusters $N$ was 20 unless otherwise specified. The maximum patience $\tau$ was varied across model sizes: it was set at 200 for DiT-S/2, DiT-B/2, and EDM, and 400 for DiT-L/2. EDM was trained using fp16, while the other models were trained using fp32. We used 8 A100 GPUs for all experiments.
Vanilla
Naive
Ours
Figure C: Qualitative comparison between vanilla, naive curriculum, and ours on the FFHQ dataset.
Vanilla
Naive
Ours
Figure D: Qualitative comparison between vanilla, naive curriculum, and ours on ImageNet dataset.
@@
Figure E: Qualitative comparison between vanilla, naive curriculum, and ours on MS-COCO dataset.
@@ Appendix F Qualitative Results
Figure C presents a qualitative analysis of the performance in unconditional facial image synthesis among the vanilla, the naive curriculum approach, and our method. Our approach demonstrates superior ability in generating realistic images.
@@
In the conditional image synthesis, we exhibit the outcomes generated by the vanilla, the naive curriculum strategy, and our proposed method. Figure D clearly shows that our methodology surpasses the competing approaches in terms of quality.
@@
To further substantiate the effectiveness of our proposed technique, we conduct a qualitative comparison of the results in the text-to-image generation task among the vanilla, the naive curriculum method, and our own approach, as depicted in Fig. E.
@@ G.1 Convergence Speed across Model Size
Figure F: We observed an increase in convergence speed across various model sizes when the proposed curriculum learning approach was applied.
By leveraging the advantages of curriculum learning in diffusion training, our method offers faster convergence than vanilla training. To further investigate this aspect, we measured FID-10K through training iterations for DiT-S and DiT-L. Figure F describes the results, showing that our curriculum approach achieves faster convergence in both models. These results also support the effectiveness of our method.
@@ G.3 Qualitative Results from DiT-L/2 with 2M iterations
In Figures G-K, we present images generated by DiT-L using our curriculum training method for 2M iterations. The results demonstrate that our method produces highly realistic images.
Figure G: Uncurated 256$\times$256 DiT-L/2 samples.
@@
Figure H: Uncurated 256$\times$256 DiT-L/2 samples.
@@
Figure I: Uncurated 256$\times$256 DiT-L/2 samples.
@@
Figure J: Uncurated 256$\times$256 DiT-L/2 samples.
@@
Figure K: Uncurated 256$\times$256 DiT-L/2 samples.
@@
Appendix H Discussion on Similarity with the Previous Work
While both our work and (Go et al., 2023a) explore the characteristics of denoising tasks in diffusion models, the aspects of exploration in each work are substantially different. The notion of task affinity introduced in (Fifty et al., 2021, Go et al., 2023a) refers to how harmoniously the model can learn multiple tasks together. Specifically, their work focuses on identifying and mitigating conflicts between tasks, emphasizing task interactions and transferability by analyzing task similarities (e.g., gradient similarity or alignment). In contrast, our work explicitly quantifies the relative difficulty of individual denoising tasks across timesteps as a standalone property, independent of task interdependencies. The analysis of task difficulty in our work involves evaluating metrics such as loss behavior or convergence rates, directly reflecting the complexity of solving each task at different timesteps. Therefore, while (Go et al., 2023a) addresses how tasks relate and interact during multi-task learning, our focus lies in systematically characterizing the intrinsic difficulty of tasks across timesteps in diffusion models.
Appendix I Broader Impacts